Find the dimensions of the right circular cone of least volume which can be

circumscribed about a right circular cylinder of radius 4 cm and height of 5 cm.


1
Expert's answer
2021-12-16T17:12:28-0500

Solution:


Given, "AB=CD=4\ cm, AC=5\ cm"

Let "AE=H\ cm, AF=R\ cm"

Then, "CE=H-5"

Since, triangle EDC and EFA are similar by AA rule, "\dfrac{EC}{EA}=\dfrac{CD}{AF}"

"\Rightarrow \dfrac{H-5}{H}=\dfrac{4}{R}
\\\Rightarrow RH-5R=4H
\\\Rightarrow H=\dfrac{5R}{R-4}\ ...(i)"


Now, volume of cone, "V=\dfrac13 \pi R^2H"

"\Rightarrow V=\dfrac13 \pi R^2(\dfrac{5R}{R-4})\ [Using \ (i)]
\\ \Rightarrow V=\dfrac53 \pi (\dfrac{R^3}{R-4})"

On differentiating w.r.t R,

"V'=\dfrac53 \pi (\dfrac{(R-4)3R^2-R^3(1)}{(R-4)^2})
\\ \Rightarrow V'=\dfrac53 \pi \dfrac{3R^3-12R^2-R^3}{(R-4)^2}
\\ \Rightarrow V'=\dfrac53 \pi \dfrac{2R^2(R-6)}{(R-4)^2}\ ...(ii)"

Now, put V'=0

"\dfrac53 \pi \dfrac{2R^2(R-6)}{(R-4)^2}=0
\\\Rightarrow R-6=0
\\\Rightarrow R=6"

Again, differentiating (ii) w.r.t. R,

"\\ V''=\dfrac{10}3 \pi \dfrac{d}{dR}(\dfrac{R^3-6R^2}{(R-4)^2})
\\=\dfrac{10}3 \pi (\dfrac{(R-4)^2(3R^2-12R)-(R^3-6R^2)2(R-4)}{(R-4)^4})
\\>0\ for \ R=6"

Thus, minima exists.

Now, put R=6 in (i).

"H=\dfrac{5(6)}{6-4}=15\ cm"

Hence, the dimensions of the right circular cone are R=6 cm, H=15 cm.